Problem Description: An ichthyologist wants to determine if a population of
Lumpsuckers is evolving. He samples 5,000 fish and determines their genotype
for an AFLP locus with two alleles: B (band) and Nb (no band). He notices that
3,500 individuals are homozygous for the B allele, with the rest of individuals
equally divided amongst heterozygotes and Nb homozygotes.
Questions:
1. How many individuals are there in the sample for the other two genotypes? (2 pts)
2. What are the observed genotypic frequencies in the sample? (3 pts)
3. What are the observed allelic frequencies in the sample? (4 pts)
4. What are the expected genotypic frequencies under HWE for the sample? (6 pts)
5. Is the population evolving? Explain (Hint: calculate the Χ2 value) (5 pts)
Answer #1

1) no.of indivisuals with heterogygous and homozygous recessive = 5000-3500=1500 which are equally divided into heterogygous indivisual which means it's no. Will be 1500/2 = 750 and homozygous recessive = 750

2) gene frequency for homozygous dominant = 3500/5000= 0.7

Gene frequency for heterozygous = 750/ 5000= 0.15

Gene frequency for homozygous recessive = 750/ 5000= 0.15

3) frequency for dominant allele=( no. Of BB + 1/2 of heterozygous)/ total individuals= (3500+1/2X750)/5000=( 3500+ 375)/5000= 3875/5000=0.775

Observed frequency for recessive allele= (no. Of homogygous recessive +1/2 of heterogygous no.)/ Total =( 750+1/2x 750)/5000= (750+375)/5000= 1125/5000=0.225

4) for Hardy Weinberg equilibrium, dominant allele = 0.5 recessive allele= 0.5

For dominant genetype frequency= (PXP)x no.of total population = 0.5x0.5x5000= 1250

For heterogygous genotype frequency= 2pqx 5000= 2x 0.5x0.5x5000= 2500

For homozygous recessive genotype frequency = qxq x 5000= 0.5x0.5x5000= 1250

5) this population is evolving as allelic frequency is shifting toward dominant type means dominant allele type is preferred as it's allelic frequency shifted from 0.5 to 0.775 , which means it's is more selected over recessive type allele .

Hope it's clear..thanks